Could someone please help me understand how to do this equation?
-8x-25=27 2/3
I know I have to multiply by three to make a whole number so 27 2/3 becomes -83
then I believe I am supposed to add either -25 or 25 to both sides

Found 2 solutions by edjones, ewatrrr:
Answer by edjones(8007)

-8x-25=27 2/3
-8x-25=83/3
-24x-75=83
-24x=158
x=-158/24=-79/12
.
Ed

Answer by ewatrrr(24785)


Hi
-8x-25 = 27 2/3 = 83/3 |27 2/3 is the same as 83/3
-8x-25 = 83/3 |multiplying thru by 3 so as all denominators = 1
-24x - 75 = 83
-24x = 158
x = -158/24 = -79/12